Dividends Per Share

Imaging Inc., a developer of radiology equipment, has stock outstanding as follows: 21,000 shares of cumulative preferred 1% stock, $110 par, and 70,000 shares of $20 par common. During its first four years of operations, the following amounts were distributed as dividends: first year, $15,540; second year, $40,660; third year, $79,240; fourth year, $153,300.

Compute the dividends per share on each class of stock for each of the four years. Round all answers to two decimal places. If no dividends are paid in a given year, enter "0".

1st Year 2nd Year 3rd Year 4th Year
Preferred stock (dividend per share) $ $ $ $
Common stock (dividend per share) $ $ $ $
Answer #1
Preferred shares 21000
Par value 110
Dividend rate 1%
Common shares 70000
Preferred stock will receive a fixed dividend every year
Preferred stock (dividend per share) 1.1
=110*1%
Preferred stock total dividend 23100
=110*1%*21000

Since the preferred shares are cumulative in nature, any dividend less than the minimum required amount as calculated above will be carried forward and paid next year. Common shares will receive dividend only when all the current as well as past (accumulated) dividend on preferred shares has been paid off

1st Year
Total dividend per share 15540
Preferred stock total dividend 15540
Preferred stock (dividend per share) 0.74 =15540/21000
Common stock (dividend per share) 0
Carried forward dividend on preferred share 0.36 =1.1-0.74
2nd Year
Total dividend per share 40660
Preferred stock total dividend 23100 =110*1%*21000
Preferred stock (dividend per share) 1.1 =110*1%
Carried forward dividend on preferred share 7560 =23100-15540
Carried forward dividend on preferred share (per share) 0.36 =7560/21000
Total dividend per share- Preferred shares 1.46 =1.1+0.36
Total dividend - Preferred shares 30660 =23100+7560
Common stock total dividend 10000 =40660-30660
Common stock (dividend per share)          0.14 =10000/70000
3rd Year
Total dividend per share 79,240
Preferred stock total dividend 23100 =110*1%*21000
Preferred stock (dividend per share) 1.1 =110*1%
Common stock total dividend 56,140 =79240-23100
Common stock (dividend per share)          0.80 =56140/70000
4th Year
Total dividend per share 153,300
Preferred stock total dividend 23100 =110*1%*21000
Preferred stock (dividend per share) 1.1 =110*1%
Common stock total dividend 130,200 =153300-23100
Common stock (dividend per share) 1.86 =130200/70000
1st Year 2nd Year 3rd Year 4th Year
Total dividend 15540 40660 79240 153300
Preferred stock (dividend per share) 0.74 1.46 1.1 1.1
Common stock (dividend per share) 0          0.14          0.80

1.86